5. A small-town carnival has a simple game that you can play with a number cube. If you roll a 

6 you win a fluffy teddy bear. If you roll a 1, 2, 3, 4, or 5 you get a piece of stale gum. After 

watching other children play the game 50 times, only 2 children have won a teddy bear. You 

decide to simulate the game with a random number table. 

a) Determine labels for your outcomes.

 

 

b) Using the random number table below, determine how many times a “6” appears in 50 trials. 

 

 

 

 

 

c) What percentage of a “6” in 50 trials did you get from the simulation? 

 


d) What percentage of a “6” did you observe at the carnival?

 



e) How do your answers for parts (c) and (d) compare? What does this tell you about the


1
Expert's answer
2021-05-20T08:29:39-0400

a:

"6^{2}=36 ,\space 5^{2}=25 , and \space 5×6=30"

"30+25+36=91"

Outcomes "6^{3}=216"


b:

Probability"=\frac {1}{216}\times 50=0.2"


c:

"0.2×100=20\%"


d: "\frac{91}{216}×50=21"


e: part c shows the probability of getting 6 in fifty trials while d shows the number of 6 one can get from the probability success.